# Free-field construction of Carrollian $W_{N}$-algebras

### v2

@@ 1 Introduction
Carrollian symmetries LevyLeblond:1965 have recently emerged as central in the exploration of non-Lorentzian limits of relativistic theories Henneaux:2021yzg , with wide-ranging applications in areas such as flat space holography Bagchi:2010zz ; Bagchi:2012cy ; Barnich:2012aw ; Duval:2014uva ; Bagchi:2016bcd ; Bagchi:2022emh ; Donnay:2022aba ; Nguyen:2023vfz ; Nguyen:2023miw ; Bagchi:2023cen , tensionless and null strings Bagchi:2013bga ; Bagchi:2015nca , black hole physics Donnay:2019jiz ; Penna:2018gfx ; Ecker:2023uwm , cosmology deBoer:2021jej ; deBoer:2023fnj , hydrodynamics Bagchi:2023ysc , the fluid/gravity correspondence Ciambelli:2018wre ; Campoleoni:2018ltl ; Campoleoni:2022wmf , and condensed matter systems Bidussi:2021nmp ; Bagchi:2022eui . The associated Carroll algebra arises as an ultra-relativistic Inönü–Wigner contraction of the Poincaré group, corresponding to the formal limit in which the speed of light is sent to zero LevyLeblond:1965 ; SenGupta:1966qer . At the level of conformal symmetries, this contraction yields the Carrollian conformal algebra, which in two dimensions coincides with the BMS3 algebra, the asymptotic symmetry algebra of three-dimensional asymptotically flat spacetimes Bondi:1962px ; Sachs:1962zza .
@@
The recognition that the BMS3 algebra is isomorphic to a contraction of two copies of the Virasoro algebra Bagchi:2010zz has provided a conceptual foundation for flat space holography, proposing that gravity in asymptotically flat spacetime admits a dual description in terms of a codimension-one Carrollian conformal field theory (CCFT) living at null infinity Bagchi:2023cen . [^1] Another approach is celestial holography Strominger:2017zoo , which provides a complementary but related perspective on flat space holography Donnay:2022aba ; Bagchi:2022emh . Besides studying supersymmetric extensions (see for example Bagchi:2009ke ; Bagchi:2022owq ; Lodato:2016alv ; Barnich:2015sca ), a natural generalization of such a holographic duality is to include higher-spin fields in the bulk. In AdS3, higher-spin gravity is known to exhibit asymptotic $W$-symmetries Campoleoni:2010zq ; Henneaux:2010xg , suggesting that its holographic duals are 2D CFTs with extended conformal symmetry generated by $W$-algebras. The flat-space analogue of this structure was investigated in Afshar:2013vka ; Gonzalez:2013oaa , where a spin-3 gravity theory in flat space was shown to possess a Carrollian version of the $W_{3}$-algebra as its asymptotic symmetry. More generally, one expects that a flat space higher-spin gravity theory (based on a corresponding flat higher-spin algebra Ammon:2017vwt ; Campoleoni:2021blr ) has an asymptotic symmetry given by the corresponding Carrollian $W$-algebra. It arises from a Carrollian contraction of the relativistic $W$-algebras [^2] Similar to the contraction of Galilean $W$-algebras Rasmussen:2017 . and extends the Carrollian conformal algebra (or BMS3 algebra) by higher-spin currents, naturally generalizing the Carrollian symmetry paradigm.
@@
which coincide with the classical expressions.
@@ 4.4 Proper quantum Carrollian $W_{N}$-algebras
$\quad=\mathcal{J}^{-}_{1}\big[\mathcal{J}^{+}_{1},\mathcal{J}^{-}_{2}\big]\mathcal{J}^{+}_{2}+\mathcal{J}^{-}_{1}\mathcal{J}^{+}_{2}\big[\mathcal{J}^{+}_{1},\mathcal{J}^{-}_{2}\big]+\mathcal{J}^{-}_{2}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{+}_{1}+{\color[rgb]{0,0,1}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{-}_{2}\mathcal{J}^{+}_{1}}$
$\qquad+{\color[rgb]{0,0,1}\mathcal{J}^{+}_{1}\mathcal{J}^{-}_{2}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]}+\mathcal{J}^{+}_{1}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{-}_{2}+\big[\mathcal{J}^{+}_{1},\mathcal{J}^{-}_{2}\big]\mathcal{J}^{+}_{2}\mathcal{J}^{-}_{1}+\mathcal{J}^{+}_{2}\big[\mathcal{J}^{+}_{1},\mathcal{J}^{-}_{2}\big]\mathcal{J}^{-}_{1}\,.$ (93)
@@
${\color[rgb]{0,0,1}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{-}_{2}\mathcal{J}^{+}_{1}+\mathcal{J}^{+}_{1}\mathcal{J}^{-}_{2}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]}=\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{+}_{1}\mathcal{J}^{-}_{2}+\mathcal{J}^{-}_{2}\mathcal{J}^{+}_{1}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big$]\,. (94)
@@ 5 Discussion and Outlook
On the other hand, the Galilean $W_{N}$-algebra can be obtained either by a Galilean contraction or by a flipped Carrollian contraction, where the time direction is reversed in one sector. As a concrete example, we worked out the full structure of the Galilean $W_{3}$-algebra using the free-field construction (see Appendix C). Whereas specific examples of free-field constructions for the Galilean conformal and the Galilean $W_{3}$-algebra have appeared in the literature Gonzalez:2014tba ; Banerjee:2015kcx ; Radobolja:2021nae , we provide a novel, systematic free-field construction for general Carrollian and Galilean $W_{N}$-algebras, which also clarifies the subtle differences between the Galilean and the Carrollian case in the quantum regime.

### v3

@@ 1 Introduction
Carrollian symmetries LevyLeblond:1965 have recently emerged as central in the exploration of non-Lorentzian limits of relativistic theories Barnich:2012rz ; Duval:2014uoa ; Henneaux:2021yzg ; deBoer:2021jej ; Rivera-Betancour:2022lkc ; Baiguera:2022lsw ; Miskovic:2023zfz , with wide-ranging applications in areas such as flat space holography Bagchi:2010zz ; Bagchi:2012cy ; Barnich:2012aw ; Duval:2014uva ; Bagchi:2016bcd ; Bagchi:2022emh ; Donnay:2022aba ; Nguyen:2023vfz ; Nguyen:2023miw ; Bagchi:2023cen , tensionless and null strings Bagchi:2013bga ; Bagchi:2015nca , black hole physics Donnay:2019jiz ; Penna:2018gfx ; Ecker:2023uwm , cosmology deBoer:2021jej ; deBoer:2023fnj , hydrodynamics deBoer:2017ing ; Ciambelli:2018xat ; Petkou:2022bmz ; Bagchi:2023ysc , the fluid/gravity correspondence Ciambelli:2018wre ; Campoleoni:2018ltl ; Campoleoni:2022wmf , and condensed matter systems Bidussi:2021nmp ; Bagchi:2022eui . The associated Carroll algebra arises as an ultra-relativistic Inönü–Wigner contraction of the Poincaré group, corresponding to the formal limit in which the speed of light is sent to zero LevyLeblond:1965 ; SenGupta:1966qer . At the level of conformal symmetries, this contraction yields the $d$-dimensional Carrollian conformal algebra, which coincides with the BMSd+1 algebra, the asymptotic symmetry algebra of $(d+1)$-dimensional asymptotically flat spacetimes Bondi:1962px ; Sachs:1962zza ; Barnich:2006av .
@@
The recognition that the BMS3 algebra is isomorphic to a contraction of two copies of the Virasoro algebra Bagchi:2010zz has provided a conceptual foundation for flat space holography, proposing that gravity in asymptotically flat spacetime admits a dual description in terms of a codimension-one Carrollian conformal field theory (CCFT) living at null infinity Bagchi:2023cen . [^1] Another approach is celestial holography Strominger:2017zoo , which provides a complementary perspective on the flat-space scattering sector Donnay:2022aba ; Bagchi:2022emh . Besides studying supersymmetric extensions (see for example Bagchi:2009ke ; Bagchi:2022owq ; Lodato:2016alv ; Barnich:2015sca ), a natural generalization of such a holographic duality is to include higher-spin fields in the bulk. In AdS3, higher-spin gravity is known to exhibit asymptotic $W$-symmetries Campoleoni:2010zq ; Henneaux:2010xg , suggesting that its holographic duals are 2D CFTs with extended conformal symmetry generated by $W$-algebras. The flat-space analogue of this structure was investigated in Afshar:2013vka ; Gonzalez:2013oaa , where a spin-3 gravity theory in flat space was shown to possess a Carrollian version of the $W_{3}$-algebra as its asymptotic symmetry. More generally, one expects that a flat space higher-spin gravity theory (based on a corresponding flat higher-spin algebra Ammon:2017vwt ; Campoleoni:2021blr ) has an asymptotic symmetry given by the corresponding Carrollian $W$-algebra. It arises from a Carrollian contraction of the relativistic $W$-algebras [^2] Similar to the contraction of Galilean $W$-algebras Rasmussen:2017 . and extends the Carrollian conformal algebra (or BMS3 algebra) by higher-spin currents, naturally generalizing the Carrollian symmetry paradigm.
@@
which coincide with the classical expressions. A schematic overview over the resulting structures can be found in figure 1.
[…]
Figure 1: Whereas Carrollian and Galilean contraction of the $W$-algebras lead to isomorphic algebras in the classical case (related by exchange of space and time), one finds two different structures for the quantum algebras. Quantum analogues of the classical isomorphism are the isomorphism between the algebra obtained by the Galilean contraction and by a flipped Carrollian contraction (where normal ordering is flipped in one sector), and, analogously the isomorphism between the Carrollian contraction and a flipped Galilean one (not discussed in this work). The two quantum algebras differ in the ordering prescription and in the structure constants, which only in the case of the quantum Galilean algebra receive quantum corrections.
@@ 4.4 Proper quantum Carrollian $W_{N}$-algebras
$\quad=\mathcal{J}^{-}_{1}\big[\mathcal{J}^{+}_{1},\mathcal{J}^{-}_{2}\big]\mathcal{J}^{+}_{2}+\mathcal{J}^{-}_{1}\mathcal{J}^{+}_{2}\big[\mathcal{J}^{+}_{1},\mathcal{J}^{-}_{2}\big]+\mathcal{J}^{-}_{2}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{+}_{1}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{-}_{2}\mathcal{J}^{+}_{1}}$
$\qquad+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{J}^{+}_{1}\mathcal{J}^{-}_{2}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]}+\mathcal{J}^{+}_{1}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{-}_{2}+\big[\mathcal{J}^{+}_{1},\mathcal{J}^{-}_{2}\big]\mathcal{J}^{+}_{2}\mathcal{J}^{-}_{1}+\mathcal{J}^{+}_{2}\big[\mathcal{J}^{+}_{1},\mathcal{J}^{-}_{2}\big]\mathcal{J}^{-}_{1}\,.$ (93)
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{-}_{2}\mathcal{J}^{+}_{1}+\mathcal{J}^{+}_{1}\mathcal{J}^{-}_{2}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]}=\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big]\mathcal{J}^{+}_{1}\mathcal{J}^{-}_{2}+\mathcal{J}^{-}_{2}\mathcal{J}^{+}_{1}\big[\mathcal{J}^{-}_{1},\mathcal{J}^{+}_{2}\big$]\,. (94)
@@ 5 Discussion and outlook
On the other hand, the Galilean $W_{N}$-algebra can be obtained either by a Galilean contraction or by a flipped Carrollian contraction, where the time direction is reversed in one sector. As a concrete example, we worked out the full structure of the Galilean $W_{3}$-algebra using the free-field construction (see Appendix C). Whereas specific examples of free-field constructions for the Galilean conformal and the Galilean $W_{3}$-algebra have appeared in the literature Gonzalez:2014tba ; Banerjee:2015kcx ; Ammon:2017vwt ; Radobolja:2021nae , we provide a novel, systematic free-field construction for general Carrollian and Galilean $W_{N}$-algebras, which also clarifies the subtle differences between the Galilean and the Carrollian case in the quantum regime.